# o Mata-Cervera0000-0001-8464-5102 on N. Vetlugin0000-0002-2480-0462 are Soci0000-0002-0149-9128 uel A. Porras0000-0001-8058-9377 ie Shen0000-0002-6700-9902

### v1

@@ Abstract
In conclusion, we have demonstrated for the first time the possibility of protecting structured light by the Zeno effect. Specifically, LG modes in random media maintain high mode purity by imparting Gouy phase kicks by simple imaging systems. The universality of the mechanisms involved —parabolic short-time decay and Gouy phase in the Schrödinger wave dynamics— suggests that the same ingredients could be adapted to protect other forms of structured waves of different nature and in other disturbing environments such as scattering media. The experimental verification of this effect may be carried out by repeatedly impinging a structured beam onto a randomly patterned spatial light modulator (SLM), mimicking the effect of an extended turbulent medium, interrupting the evolution with $2f$ or $4f$ imaging systems followed by analysis of the OAM spectrum with another SLM. Our proposal should also work to protect the purity of the quantum eigenstate of OAM of single photons. In this respect, further research would point towards the protection of entangled photons with OAM by the optical Zeno effect.
Acknowledgments. The authors acknowledge no conflict of interest.
Author contributions.
Funding.
@@
I Annex
I.1 Laguerre-Gaussian modes
@@
I.2 Simulation of Kolmogorov’s turbulence
The statistics of turbulent media were analyzed by Kolmogorov using structure functions, which describe the mean fluctuations of the thermodynamic variables such as pressure or temperature [kolmogorov1991local, turb_mask_1, turb_mask_2]. These fluctuations induce spatial and temporal variations of the optical properties of the atmosphere. In particular, the random refractive index of a turbulent atmosphere can be described by a power spectral density given by Von Kármán
@@ I.2 Simulation of Kolmogorov’s turbulence
where $C_{n}^{2}$ is the refractive index structure constant, a measure of the strength of the refractive index fluctuations, $\kappa_{m}=5.92/l_{0}$, $\kappa_{0}=2\pi/L_{0}$ with $l_{0}$ and $L_{0}$ the inner and outer scales of turbulence. $\bm{\kappa}$ is the coordinate vector in reciprocal space with $\kappa^{2}=\kappa_{x}^{2}+\kappa_{y}^{2}+\kappa_{z}^{2}$. In our simulations we are in the inertial sub-range $l_{0}\ll r\ll L_{0}$, in which air flow is isotropic and Kolmogorov’s theory remains valid. A key magnitude that determines the optical disturbance introduced by turbulence is the Fried parameter. It has dimensions of length, and represents the distance between two points of the medium at which the refractive indices become uncorrelated [structlightturb], also known as coherence length
@@
where we considered constant $C_{n}^{2}$ in the propagation [structlightturb]. Here, $k=2\pi/\lambda$ is the wavenumber, $z_{\rm max}$ is the length of the turbulent medium.
The refractive index of turbulence $\Delta n(x,y,z)$ is generated in a volume $L_{x}\times L_{y}\times z_{\rm max}$, where $L_{x}=L_{y}$ are the transverse dimensions of the simulation window and $z_{\rm max}$ the total propagation distance. To construct it we follow a standard Fourier transform technique [turb_mask_1, turb_mask_2], where the refractive index Fourier coefficients $c_{\kappa_{x},\kappa_{y},\kappa_{z}}$ are Gaussian random variables with zero mean $\langle c_{\kappa_{x},\kappa_{y},\kappa_{z}}\rangle=0$ and variance given by Eq. (B1) $\langle|c_{\kappa_{x},\kappa_{y},\kappa_{z}}|^{2}\rangle=\Phi_{n}(\kappa)$. Then, the refractive index in real space $\Delta n(x,y,z)$ is obtained by 3D inverse Fourier transform with the addition of $3^{3}=27$ sub-harmonics to increase the numerical accuracy [imaging_turbulence].
We solve Eq. (1) with a split-step approach [turb_mask_1, turb_mask_2] by discretizing a refractive index volume into $\mathcal{P}$ discrete propagation planes within the total propagation distance $z_{\rm max}$. Numerically, between two consecutive planes the wave evolves a free-space distance $\delta=z_{\rm max}/\mathcal{P}$ followed by a phase mask extracted from the volume refractive index distribution at that plane [taha1984analytical].
I.3 Dimensional analysis
Two different evolutions in turbulent media share the same dynamics when the associated adimensional variables match [klug2023robust]. The disturbance experienced by the beam is determined by the adimensional variable $r_{0}/w_{0}$, where $r_{0}$ is the coherence length (B2) and $w_{0}$ is the beam width. Here we also impose that two beams experience the same diffraction, which in the LG beam optics implies that the dimensionless distance $z_{\rm max}/z_{R}$ is the same. Here $z_{R}$ is the Rayleigh range, and $z_{\rm max}$ is the total propagation distance. Between two evolutions within $z_{\rm max}$ and $z^{\prime}_{\rm max}$ we get the following equality

### v2

@@ V Conclusions
In conclusion, we have demonstrated for the first time the possibility of protecting structured light by the Zeno effect. Specifically, LG modes in random media maintain high mode purity by imparting frequent Gouy phase kicks by simple imaging systems. The universality of the mechanisms involved —parabolic short-time decay and Gouy phase in the Schrödinger wave dynamics— suggests that the same ingredients could be adapted to protect other forms of structured waves of different nature and in other disturbing environments such as scattering media. The experimental verification of this effect may be carried out by repeatedly impinging a structured beam onto a randomly patterned spatial light modulator (SLM), mimicking the effect of an extended turbulent medium, interrupting the evolution with $2f$ or $4f$ imaging systems followed by analysis of the OAM spectrum with another SLM. Our proposal should also work to protect the purity of the quantum eigenstate of OAM of single photons. In this respect, further research would point towards the protection of entangled photons with OAM by the optical Zeno effect.
Acknowledgments
@@
Data availability
The simulation data is openly available at this link.
VI Appendix
VI.1 Laguerre-Gaussian modes
@@
VI.2 Simulation of Kolmogorov’s turbulence
The statistics of turbulent media were analyzed by Kolmogorov using structure functions, which describe the mean fluctuations of the thermodynamic variables such as pressure or temperature [35, 46, 36]. These fluctuations induce spatial and temporal variations of the optical properties of the atmosphere. In particular, the random refractive index of a turbulent atmosphere can be described by a power spectral density given by Von Kármán
@@ VI.2 Simulation of Kolmogorov’s turbulence
where $C_{n}^{2}$ is the refractive index structure constant, a measure of the strength of the refractive index fluctuations, $\kappa_{m}=5.92/l_{0}$, $\kappa_{0}=2\pi/L_{0}$ with $l_{0}$ and $L_{0}$ the inner and outer scales of turbulence. $\bm{\kappa}$ is the coordinate vector in reciprocal space with $\kappa^{2}=\kappa_{x}^{2}+\kappa_{y}^{2}+\kappa_{z}^{2}$. In our simulations we are in the inertial sub-range $l_{0}\ll r\ll L_{0}$, in which air flow is isotropic and Kolmogorov’s theory remains valid. A key magnitude that determines the optical disturbance introduced by turbulence is the Fried parameter. It has dimensions of length, and represents the distance between two points of the medium at which the refractive indices become uncorrelated [10], also known as coherence length
@@
where we considered constant $C_{n}^{2}$ in the propagation [10]. Here, $k=2\pi/\lambda$ is the wavenumber, $z_{\rm max}$ is the length of the turbulent medium.
The refractive index of turbulence $\Delta n(x,y,z)$ is generated in a volume $L_{x}\times L_{y}\times z_{\rm max}$, where $L_{x}=L_{y}$ are the transverse dimensions of the simulation window and $z_{\rm max}$ the total propagation distance. To construct it we follow a standard Fourier transform technique [46, 36], where the refractive index Fourier coefficients $c_{\kappa_{x},\kappa_{y},\kappa_{z}}$ are Gaussian random variables with zero mean $\langle c_{\kappa_{x},\kappa_{y},\kappa_{z}}\rangle=0$ and variance given by Eq. (B1) $\langle|c_{\kappa_{x},\kappa_{y},\kappa_{z}}|^{2}\rangle=\Phi_{n}(\kappa)$. Then, the refractive index in real space $\Delta n(x,y,z)$ is obtained by 3D inverse Fourier transform with the addition of $3^{3}=27$ sub-harmonics to increase the numerical accuracy [47].
We solve Eq. (1) with a split-step approach [46, 36] by discretizing a refractive index volume into $\mathcal{P}$ discrete propagation planes within the total propagation distance $z_{\rm max}$. Numerically, between two consecutive planes the wave evolves a free-space distance $\delta=z_{\rm max}/\mathcal{P}$ followed by a phase mask extracted from the volume refractive index distribution at that plane [50].
VI.3 Dimensional analysis
Two different evolutions in turbulent media share the same dynamics when the associated adimensional variables match [34]. The disturbance experienced by the beam is determined by the adimensional variable $r_{0}/w_{0}$, where $r_{0}$ is the coherence length (B2) and $w_{0}$ is the beam width. Here we also impose that two beams experience the same diffraction, which in the LG beam optics implies that the dimensionless distance $z_{\rm max}/z_{R}$ is the same. Here $z_{R}$ is the Rayleigh range, and $z_{\rm max}$ is the total propagation distance. Between two evolutions within $z_{\rm max}$ and $z^{\prime}_{\rm max}$ we get the following equality